OpenStudy (anonymous):

What is the slope of the line 5x - 2y = 10?

OpenStudy (anonymous):

the slope of equation given in the form ax+by=c is given by m=-a/b here a=5 ,b=-2 can you find now ?

OpenStudy (gw2011):

The slope is: y=mx+b where m = slope 2y=5x-10 y=(5/2)x-2 m=slope=5/2

OpenStudy (anonymous):

thank you!

OpenStudy (gw2011):

Sorry, the equation should be y=(5/2)x-5 I mistyped the -10

OpenStudy (anonymous):

oohh ok :D